Ive been getting stuck on 4(3t+4)-20=3+5t solve for T
Answer by nerdybill(2448)

Isolate the unknown (t) to one side of the equation:
4(3t+4)-20=3+5t
Distribute the 4 to terms inside parenthesis:
12t+16-20=3+5t
Combine like-terms:
12t-4=3+5t
Subtracting 5t from both sides:
7t-4=3
Adding 4 to both sides:
7t=7
Finally, divide both sides by 7:
t = 1